Determine whether if lim f(c) = f(c)


x + c



1. f(x) = x+2; c = -1



2. f(x) = x-2; c = 0



3. (at c = -1 )


f(x) = {x ² - 1 if x = < -1}


f(x) = { (x - 1) ² - 4 if x = ≥ - 1}



4. (at c = 1 )


f(x) = {x³ - 1 if x = < 1}


f(x) = { x² +4 if x = ≥ 1}

1
Expert's answer
2022-03-04T12:25:17-0500

1, The function "f(x)=x+2" is continuous on "\R" as a polynomial.

Therefore


"\lim\limits_{x\to c}f(x)=f(c), c\in \R"

Then


"\lim\limits_{x\to -1}f(x)=f(-1)=-1+2=1"


2, The function "f(x)=x-2" is continuous on "\R" as a polynomial.

Therefore


"\lim\limits_{x\to c}f(x)=f(c), c\in \R"

Then


"\lim\limits_{x\to0}f(x)=f(0)=0-2=-2"



3.


"\lim\limits_{x\to -1^-}f(x)=(-1)^2-1=0"

"\lim\limits_{x\to -1^+}f(x)=(-1-1)^2-4=0"

Since


"\lim\limits_{x\to -1^-}f(x)=0=\lim\limits_{x\to -1^+}f(x),"

then "\lim\limits_{x\to -1}f(x)" exists, and "\lim\limits_{x\to -1}f(x)=0."


4.


"\lim\limits_{x\to 1^-}f(x)=(1)^3-1=0"

"\lim\limits_{x\to 1^+}f(x)=(1)^2+4=5"

Since


"\lim\limits_{x\to 1^-}f(x)=0\not=5=\lim\limits_{x\to 1^+}f(x),"

then "\lim\limits_{x\to 1}f(x)" does not exist.